In a survey of 300 random households in a particular village, mother's have been asked if they used cellphone to communicate there are 76 who said YES, Use 90%, 95% and 99% confidence to estimate the proportion of all mothers who use the cellphone to communicate

1
Expert's answer
2022-05-02T16:33:28-0400

The sample proportion is computed as follows, based on the sample size "N = 300" and the number of favorable cases "X = 76:"

"\hat{p}=\dfrac{X}{N}=\dfrac{76}{300}\approx0.253333"

1. The critical value for "\alpha = 0.10" is "z_c = z_{1-\alpha/2} = 1.6449."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(\dfrac{76}{300}-1.6449\sqrt{\dfrac{\dfrac{76}{300}(1-\dfrac{76}{300})}{300}},"

"\dfrac{76}{300}+1.6449\sqrt{\dfrac{\dfrac{76}{300}(1-\dfrac{76}{300})}{300}})"

"=(0.212,0.295)"

Therefore, based on the data provided, the 90% confidence interval for the population proportion is "0.212 < p < 0.295," which indicates that we are 90% confident that the true population proportion "p" is contained by the interval "(0.212, 0.295)."


2. The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(\dfrac{76}{300}-1.96\sqrt{\dfrac{\dfrac{76}{300}(1-\dfrac{76}{300})}{300}},"

"\dfrac{76}{300}+1.96\sqrt{\dfrac{\dfrac{76}{300}(1-\dfrac{76}{300})}{300}})"

"=(0.204,0.303)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.204 < p < 0.303," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.204, 0.303)."


3. The critical value for "\alpha = 0.01" is "z_c = z_{1-\alpha/2} = 2.5758."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(\dfrac{76}{300}-2.5758\sqrt{\dfrac{\dfrac{76}{300}(1-\dfrac{76}{300})}{300}},"

"\dfrac{76}{300}+2.5758\sqrt{\dfrac{\dfrac{76}{300}(1-\dfrac{76}{300})}{300}})"

"=(0.189,0.318)"

Therefore, based on the data provided, the 99% confidence interval for the population proportion is "0.189< p < 0.318," which indicates that we are 99% confident that the true population proportion "p" is contained by the interval "(0.189, 0.318)."